OpenStudy (anonymous):

A child's sled, initially at rest, is pushed while on ice with a constant horizontal force of 20.0 N. The total mass of sled plus cargo is 20.0 kg. How fast is the sled moving after 2.00 s of being pushed? Ignore friction. A) 1.00 m/s B) 2.00 m/s C) 9.80 m/s D) 20.0 m/s

OpenStudy (azureilai):

F=ma 20N/20kg=a

OpenStudy (anonymous):

so A is the answer? @Azureilai

OpenStudy (anonymous):

No it isnt

OpenStudy (azureilai):

but it should be, it is only asking about applied force

OpenStudy (anonymous):

It's asking about final velocity

OpenStudy (anonymous):

Which is why it has a time component included

OpenStudy (anonymous):

@KeithAfasCalcLover so how do i find the answer for this?

OpenStudy (anonymous):

Ill write it out for ya :)

OpenStudy (anonymous):

So we know that: \[\eqalign{ &F=20.0N \\ &m=20.0kg \\ &v_i=0m/s \\ &\Delta t=2.00s }\] So then we know that: \[F=ma\rightarrow a=\frac{F}{m}\] We also know that: \[a=\frac{\Delta v}{\Delta t}=\frac{v_f-v_i}{\Delta t}\rightarrow v_f=a\Delta t+v_i\] Since we know \(a=\frac{F}{m}\), we can substitute: \[v_f=\left(\frac{F}{m}\right)\Delta t+v_i\] And we obtain: \[v_f=\frac{F\Delta t}{m}+v_i\] We can plug in what we know: \[v_f=\frac{20.0(2.00)}{20.0}+0=\frac{40.0}{20.0}=2.00m/s\] In conclusion, Since \(B.\phantom{..}2.00m/s\), therefore, the answer is B

OpenStudy (azureilai):

Oh ok, I see what you are talking about.